# Generalizable Articulated Object Reconstruction from Casually Captured RGBD Videos

### v1

@@ Abstract
Articulated objects are prevalent in daily life. Understanding their kinematic structure and reconstructing them have numerous applications in embodied AI and robotics. However, current methods require carefully captured data for training or inference, preventing practical, scalable, and generalizable reconstruction of articulated objects. We focus on reconstruction of an articulated object from a casually captured RGBD video shot with a hand-held camera. A casually captured video of an interaction with an articulated object is easy to acquire at scale using smartphones. However, this setting is quite challenging, as the object and camera move simultaneously and there are significant occlusions as the person interacts with the object. To tackle these challenges, we introduce a coarse-to-fine framework that infers joint parameters and segments movable parts of the object from a dynamic RGBD video. To evaluate our method under this new setting, we build a 20 larger synthetic dataset of 784 videos containing 284 objects across 11 categories. We compare our approach with existing methods that also take video as input. Experiments show that our method can reconstruct synthetic and real articulated objects across different categories from dynamic RGBD videos, outperforming existing methods significantly.
@@ 1 Introduction
Articulated objects are prevalent in daily life. Building digital representations of such objects and understanding their kinematic structures has many applications in computer animation, embodied AI, and robotics. For example, automatically building interactable digital twins of articulated objects can benefit robotics research, since researchers can train and test manipulation policies more conveniently and efficiently in simulation prior to real-world deployment. However, to the best of our knowledge, current approaches to articulated object reconstruction impose many constraints on the input data, preventing practical and efficient deployment for acquisition of articulated objects in realistic settings. Thus, researchers still largely rely on manual annotation of articulation joint parameters [39].
@@
Recent approaches to reconstructing articulated objects can be categorized into three families. The first uses a deformation model for implicit functions to represent the dynamics of articulated objects [26, 40, 46, 34]. However, the resulting implicit representation is not easy to use in simulation as we first need to reconstruct the articulated object explicitly and find rigid transformations of the different object parts. Thus, this is not a practical path for physically accurate animation or interaction. The second family reconstructs articulated objects by estimating joint parameters explicitly from two different states of the object [8, 16, 47, 19, 21, 27]. However, this type of approach requires aligning observations of two states to the same coordinate frame. Prediction accuracy can drop drastically due to slight errors or perturbations in the coordinate frame alignment of observations at each state. The third family of approaches generates articulated objects leveraging large language models and vision foundation models to help predict joint parameters [22, 13]. At present, these approaches either rely on an external library for object retrieval or require additional data for fine-tuning. Retrieving objects from a library is not easily scalable, while collecting and annotating data for fine-tuning is labor-intensive. Furthermore, these approaches do not easily generalize well to real settings after fine-tuning on limited data.
@@
By analyzing the issues of recent approaches, we summarize three desirable features for a practical articulated object reconstruction pipeline: 1) minimal constraints on input; 2) explicit joint parameter and geometry representation for easy integration into simulators; and 3) generalization across diverse categories of articulated objects. To address these desiderata, we propose to reconstruct articulated objects from casually captured RGBD videos of interaction with the object (i.e. a video shot from a hand-held RGBD camera while the same person recording interacts with the object). Casually captured videos, or dynamic videos, are easy to capture at scale and contain the necessary motion information to infer joint parameters. Compared to the problem setting in prior work, the input video data is far less constrained and easier to capture, the resulting explicit representation of articulated objects is easier to use for simulation, and the overall pipeline is scalable to real-world settings.
@@
However, this kind of input also brings more challenges to this problem. Since both the camera and the scene are dynamic, and the person occludes the object as they are manipulating it, decoupling camera motion and object part motion is challenging. Consequently, accurate estimation of joint parameters and movable part segmentation is difficult. To address these challenges, we develop a coarse-to-fine framework to reconstruct articulated objects from dynamic RGBD videos, illustrated in Fig. 1. Our pipeline first estimates joint parameters and a movable part segmentation via image feature matching. Then, a gradient-based optimization framework refines these initial estimates against a surface point cloud representation of the object acquired through 3D reconstruction. Our pipeline explicitly parameterizes the articulation joint parameters and other relevant parameters, satisfying the second feature. Moreover, to aid generalization and satisfy the third desideratum, our pipeline neither relies on an external library nor requires additional data for fine-tuning. Instead, it only uses pretrained models.
@@
Since the problem setting and task setup we described has not been previously addressed, we build a new dataset for evaluation. We select 284 synthetic objects from 11 categories in the PartNet-Mobility dataset [25] to generate dynamic videos for input and evaluate the performance of our method. Compared to datasets used in prior works such as PARIS [16] and ArtGS [19], our dataset contains 20 more objects and thus provides a more robust evaluation of the performance of different methods for reconstructing articulated objects from dynamic videos.
In summary, we make the following contributions:
@@
We address the task of articulated object reconstruction from a casually captured RGBD video, a new problem setting that is more practical than prior work.
@@
To tackle this new problem, we develop a coarse-to-fine framework that reconstructs articulated objects explicitly and does not rely on an external articulated object library or additional data for fine-tuning.
@@
To evaluate our method robustly, we build a new synthetic dataset with 20 more objects than prior datasets. Experiments show that our method outperforms state of the art methods on reconstruction of both synthetic and real articulated objects across different categories.
@@ 2 Related work
Our work lies at the intersection of articulated object reconstruction and dynamic scene understanding. We review recent work in both fields to provide a comprehensive background of our work.
@@ 2.1 Articulated object reconstruction
Articulated object reconstruction is a long-standing task in computer vision and robotics. One line of research focuses on the visual fidelity of the dynamics of the articulated objects [26, 40, 46, 34]. They usually apply implicit representations and deformation fields to model the articulated objects. This type of representation is not suitable for physically accurate animation and interaction. It requires extra effort to build an explicit representation to integrate with the simulator. In this paper, we aim to build a tool for reconstructing an interactable digital twin of articulated objects, where explicitly modeling the object and inferring joint parameters is essential. Thus, this line of work is not suitable for our problem setting.
Another line of work attempts to infer joint parameters and build a digital twin of articulated objects from multi-view observations of different object states [8, 16, 47, 19]. These works implicitly assume all the observations have already been aligned to the same coordinate. In practice, human operators need to manually align the camera coordinates of all the observations, which can not scale to large amounts of data. Similarly, in robotics community, researchers use robots to interact with the articulated object and reconstruct articulated objects from the interaction [21, 27, 51, 45]. Since robots usually stay still while manipulating objects, different observations at different object states can be easily aligned with the robot’s coordinate. However, robots are expensive and may damage themselves during the interaction. Therefore, it’s not an efficient way to build a digital twin of articulated objects. Robot-See-Robot-Do [10] recovers 3D motions of the articulated object from a monocular human demonstration video. Their problem setting is close to us but requires a static view video and is sensitive to the camera view.
Recently, with the advancement of large language models and generative models, some researchers try to leverage those models to directly generate code representation or kinematic graph from a single image of an articulated object [22, 13, 2, 17]. With the power of vision-language models, Articulate-Anything can also take text or video as input [13]. These works either rely on existing mesh libraries to retrieve object meshes or require extra data for fine-tuning foundation models. Thus, at the current stage, this line of work still suffers from the out-of-distribution issue, which prevents it from being fully generalizable and scalable. In contrast, our method does not rely on object retrieval or fine-tuning, and is inherently more scalable.
@@ 2.2 Dynamic scene reconstruction
Dynamic scene understanding and reconstruction is a challenging task in computer vision and is still an open problem. One line of research continues extending neural radiance field and 3D gaussian splatting to represent dynamic scenes [29, 3, 54, 20, 48, 41, 50, 43]. These works achieve impressive results in the novel view rendering of dynamic scenes while focusing less on geometry reconstruction.
Another line of approaches attempts to reconstruct dynamic scenes from monocular depth estimation and optical flow. Robust-CVD proposes to first estimate the depth of each frame and then optimize camera poses and depth jointly [12]. CasualSAM leverages optical flow to train a depth and movement prediction network and optimize camera parameters [56]. MegaSaM proposes an optimization framework for dynamic scene reconstruction based on DROID-SLAM [14, 38]. Recently, DUSt3R [44] provides a powerful foundation model for static scene reconstruction in point cloud representation. Based on DUSt3R, better methods come out recently. MonST3R fine-tunes DUSt3R on dynamic scene datasets to recover dynamic 3D scenes [55]. CUT3R introduces a continuous 3D perception model that can perform online dense 3D reconstruction [42]. Easi3R identifies moving objects in the video from DUSt3R’s attention maps [1]. Aether proposes a unified model for reconstruction, planning, and prediction on dynamic scene [36]. Our coarse prediction pipeline directly benefits from MonST3R [55], and our refinement framework is heavily inspired by the optimization framework in CasualSAM [56] and Mega-SaM [14].
Another important task in dynamic scene understanding is segmenting the moving map, images indicating which pixels are moving in the scene at each frame, in the video. Recent approaches are mainly based on optical flow or point tracking [52, 24, 23, 9, 5, 7]. However, we find that optical flow or point tracking-based methods struggle with textureless objects and are sensitive to the points’ location. Therefore, we don’t use them in our work. We further discuss this in Sec. C.5.
@@ 3 Method
In this section, we first state the problem formulation. Then, we introduce our coarse-to-fine reconstruction pipeline in the next two subsections.
@@ 3.1 Problem formulation
In this paper, we proposed to reconstruct articulated object from a casually capture RGBD video , where denotes the video length and each frame represent an RGBD image with height and width . To achieve accurate reconstruction of an articulated object, it is essential to estimate two key components: 1) The joint parameters , including the type of joint (e.g., revolute or prismatic), as well as the position and orientation of the joint axis. 2) The movable parts of the object. Since it’s challenging to segment movable parts on the 3D objects directly, we instead estimate the moving map of each video frame , which indicates the moving parts of the object in the video. Then, we reproject the moving map to 3D space with depth values to identify movable parts of the object.
For the input video , we assume it is causally captured using handheld devices such as mobile phones. Therefore, we do not assume that the camera remains stationary during recording, nor do we assume that the camera trajectory can be easily obtained. Both camera pose and the joint state at each frame change simultaneously, which also needs to be estimated during the reconstruction. Normally, a human operator manipulates one joint at a time. Thus, we assume only one part of the object is moving in the video.
@@ 3.2 Coarse prediction
To address this issue, we first design a coarse prediction pipeline as shown in Fig. 2. The idea is to decouple the problem by firstly finding the moving part in the video . Once we identify the dynamic region and static region in the video frames, we can use point motion in the static region to estimate the camera pose at each frame. We align all the observations to the same coordinate, and compute rigid transformations of movable parts of the object using points motion in the dynamic region to estimate joint parameters and joint state .
In practice, we use MonST3R [55] to provide a rough moving map of each video frame. Then we reproject pixels in to 3D point cloud with camera intrinsics, and utilize LoFTr [35] to compute pixel-level feature matching between two video frames. We use feature matching in static regions identified by to compute relative camera pose. Then, we align all observations to the same camera coordinate with the relative camera poses of each frame. Then, we use feature matching in the dynamic regions to compute a rigid transformation of the movable parts of the object. We can further estimate joint parameters from this rigid transformation. We estimate joint parameters from multiple video frame pairs and average the results as the coarse prediction of the joint parameters and states.
Through the coarse prediction stage, the proposed pipeline obtains rough results on moving map segmentation and joint parameters. However, this method has two disadvantages: 1) Since the moving map is inaccurate, camera poses and joint parameters estimation are inaccurate. 2) Feature matching provides a reliable prediction only when two video frames are visually similar. In other words, the joint states do not change much, and the point motion in dynamic regions will be minimal. In this situation, it’s hard to tell which joint types lead to this kind of point motion. Joint type prediction will be inaccurate as well.
@@ 3.3 Refinement
We further refine the estimation results through gradient-based optimization. It is based on a simple idea: if the estimation of the joint parameters and camera poses is accurate, the object states observed in the video should be able to transform back perfectly to the initial state. In other words, the transformed point cloud should align with the given object surface point cloud at the initial state. An overview of our pipeline is shown in Fig. 3.
@@
Concretely, we divide the optimization target into static part and moving part. For the static part, we transform the point cloud to the same coordinate with camera poses . We compute the single directional chamfer distance to the given object surface point cloud in Eq. 1.
@@
To compute the chamfer distance of the moving part, we first compute the transformation matrix for the moving part with and . Then, we transform with camera pose and . We compute the single directional chamfer distance to in Eq. 2.
@@
Combining these two equations produces the final optimization objective . Note that this computation graph is differentiable. Thus, we can simply use gradient-based optimization methods to optimize joint parameters, camera poses, and the moving map.
@@
However, this implementation has two problems: 1) how to deal with the newly observed part of the object in the video. For example, the interior of a microwave cannot be observed when the door is closed at the beginning. However, it will be observed during the interaction with the door. This interior part does not correspond with the object surface point cloud. Consequently, chamfer distance will provide an inaccurate estimation of point cloud alignment, resulting in biased gradients. 2) how to optimize efficiently. In the previous formulation, we optimize the probability of each pixel belonging to the moving part of each frame, which contains parameters to optimize.
The above two problems share a similar property: both of them care about part-level information. The first one cares about whether a part is newly observed, while the second one cares about which parts are moving. If we can automatically segment the whole video into different parts and track them, we can focus on which parts are newly observed and which are moving. In this case, the number of parts is much smaller than the number of pixels, reducing the optimization complexity.
In practice, we first find the parts that appear in the initial frame and then filter out parts in the following frames that do not appear in the initial frame. This resolves the false correspondence issue when computing the chamfer distance. For the second problem, instead of optimizing on each pixel, we optimize of each segmented part. We define a vector to represent the probability of each part being a moving part, where is the number of segmented parts. Then, for each part , we can compute the probability of being a moving part by multiplying and the mask of this part . We can construct simply by aggregating the mask of each part . We adopt an existing implementation of automatic part segmentation on videos. Finally, we find movable parts on via finding points in the dynamic regions in . We adopt NKSR [6] to reconstruct meshes.

### v2

# iTACO: Interactable Digital Twins of Articulated Objects from Casually Captured RGBD Videos

@@ Abstract
Articulated objects are prevalent in daily life. Interactable digital twins of such objects have numerous applications in embodied AI and robotics. Unfortunately, current methods to digitize articulated real-world objects require carefully captured data, preventing practical, scalable, and generalizable acquisition. We focus on motion analysis and part-level segmentation of an articulated object from a casually captured RGBD video shot with a hand-held camera. A casually captured video of an interaction with an articulated object is easy to obtain at scale using smartphones. However, this setting is challenging due to simultaneous object and camera motion and significant occlusions as the person interacts with the object. To tackle these challenges, we introduce iTACO: a coarse-to-fine framework that infers joint parameters and segments movable parts of the object from a dynamic RGBD video. To evaluate our method under this new setting, we build a dataset of 784 videos containing 284 objects across 11 categories that is 20 larger than available in prior work. We then compare our approach with existing methods that also take video as input. Our experiments show that iTACO outperforms existing articulated object digital twin methods on both synthetic and real casually captured RGBD videos.
@@ 1 Introduction
Building digital twins of articulated objects has many applications in computer animation, embodied AI, and robotics. For example, scalable pipelines for building interactable digital twins benefit robotics research, by enabling convenient training and evaluation in simulation prior to real-world deployment. However, current approaches to building interactable replicas of articulated objects impose many constraints on the input data, preventing practical and efficient acquisition of articulated objects in realistic settings. Thus, researchers still largely rely on manual annotation of articulation joint parameters [53].
@@
Recent approaches to articulated object reconstruction and motion analysis can be categorized into three families. The first focuses on dynamic reconstruction of the shape of articulated objects [34, 78, 54, 62, 44]. Unfortunately, the final output is not interactable as it is typically a sequence of meshes instead of a kinematic structure with explicit links and joints. The second family infers joint parameters and builds part-level geometry simultaneously from two different states of the object [12, 24, 64, 27, 29, 36]. However, this requires aligning observations between states, and even small errors in alignment can lead to dramatic drops in accuracy. The third family of approaches generates articulated objects leveraging large language and multimodal models [30, 18]. At present, these approaches either rely on an external library for object retrieval or require data for fine-tuning. Retrieving objects is not easily scalable, while collecting data for fine-tuning is labor-intensive. Furthermore, generalization to real settings after fine-tuning on limited data is not easy.
@@
By analyzing the issues of recent approaches, we summarize three desirable features for practical interactable digital twin acquisition: 1) minimal constraints on input; 2) explicit joint parameter and part-level geometry representation for easy integration into simulators; and 3) generalization across diverse object categories. To address these desiderata, we create digital twins of articulated objects from casually captured RGBD videos showing a human interacting with the object. For example, a human operator with a hand-held camera can first scan the object to reconstruct the object surface, and then record an interaction with the object. Such videos are easy to capture at scale and contain the necessary motion information to infer joint parameters. Compared to the problem setting in prior work, casually captured videos are far less constrained and easier to capture, the final output is easier to use for simulation, and the overall pipeline is scalable to real-world settings.
@@
However, casually-captured video also brings more challenges. Since both the camera and the scene are dynamic, and the person occludes the object as they are manipulating it, decoupling camera motion and object part motion is challenging. Consequently, accurate estimation of joint parameters and movable part segmentation is difficult. To address these challenges, we develop iTACO: a coarse-to-fine framework to build digital twins of articulated objects from dynamic RGBD videos, illustrated in Fig.˜1. Our pipeline first estimates joint parameters and a movable part segmentation via image feature matching. Then, a gradient-based optimization framework refines the initial estimates against a surface point cloud representation of the object acquired through 3D reconstruction. Our pipeline explicitly parameterizes articulation and geometric parameters, satisfying the second desideratum. Moreover, to aid generalization and satisfy the third desideratum, our pipeline neither relies on an external library nor requires additional data for fine-tuning. Instead, it only uses pretrained models.
@@
Since our problem setting has not been previously addressed, we build a dataset for evaluation. We generate videos for 284 synthetic objects from 11 categories in PartNet-Mobility [33] and collect real casual RGBD videos. Our dataset contains 20 more objects compared to prior works [24, 27], and allows a more robust evaluation of articulated object acquisition. In summary, we make the following contributions:
@@
We address building interactable digital twins of articulated objects from casually captured RGBD videos, a new problem setting that is more practical than prior work.
@@
We develop iTACO: a coarse-to-fine framework for acquisition of digital twins of articulated objects.
@@
We collect a dataset with 20 more objects than prior datasets. Experiments show that iTACO outperforms state-of-the-art methods on both synthetic and real articulated objects across different categories.
@@ 2 Related Work
Our work is related to motion analysis and reconstruction of articulated objects, and dynamic scene understanding. We review recent work in both fields.
@@ 2.1 Articulated Object Reconstruction
Motion analysis and reconstruction of articulated objects is a long-standing task in computer vision and robotics. Previous motion analysis works focus on predicting joint parameters or part-level segmentation from an image or a single point cloud [60, 74, 20, 63, 11, 48, 81]. These works require training on specific datasets, which limits generalization. Previous articulated object reconstruction works focus on reconstructing the shape of articulated objects [34, 77, 78, 62, 44]. The final output of these methods is the entire shape of the object, which is not interactable, in particular due to the absence of a kinematic structure with joints and articulation parameters that can be driven kinematically or through physics-based simulation. In contrast, we aim to build interactable digital twins of articulated objects, where inferring joint parameters and part-level geometry is essential.
Another line of work attempts to infer joint parameters and build part-level geometry of articulated objects simultaneously. One problem setting uses multi-view observations of different object states [12, 24, 64, 27]. These works implicitly assume all observations have been aligned to the same coordinates. In practice, this requires either manual alignment which is not scalable or algorithmic alignment which is error-prone. Other prior work uses robots to interact with the articulated object and reconstruct articulated objects from the interaction [29, 36, 75, 61]. However, robots are expensive and may damage themselves or the environment during the interaction. Therefore, this is not a practically scalable path to build digital twins of articulated objects. Recently, researchers began exploring motion analysis from videos [14, 65, 85]. Robot-See-Robot-Do [14] recovers 3D motions of an articulated object from a monocular human demonstration video. Their problem setting is close to us, but requires the camera to be static.
Other work uses large language and multimodal models to generate a code representation or kinematic graph from a single image of an articulated object [30, 18, 4, 25]. Leveraging vision-language models, Articulate-Anything can also take image or video as input [18]. These works either rely on existing mesh libraries to retrieve object part meshes or require extra data for fine-tuning foundation models. Thus, at the current stage, this line of work still suffers from out-of-distribution issues, limiting generalization and scalability. In contrast, our method does not rely on object retrieval or fine-tuning, and is inherently more general and scalable.
In computer animation, automatic skeletal rigging has been studied for decades [49, 1, 72, 55, 73, 70, 19, 71, 47, 45, 5, 23, 83]. However, the skeleton model is designed to animate humans and animals, whose links are non-rigid and joints are mainly ball joints. For rigid articulated objects such as drawers and microwaves, most joints are prismatic joints and revolute joints. Thus, it is suitable to represent rigid articulated objects using skeleton models.
@@ 2.2 Dynamic Scene Reconstruction
Dynamic scene reconstruction is a challenging task in computer vision. One line of research describes a dynamic scene with a deformation field. Signed distance functions, and later neural radiance fields and 3D gaussian splatting are commonly used [35, 2, 43, 38, 6, 80, 28, 66, 56, 69, 59]. These works achieve impressive results in shape reconstruction and novel view rendering of dynamic scenes, but it is difficult to transform the deformation field to an explicit motion representation, such as a rigid transformation matrix, which is necessary for our problem.
Another line of approaches attempts to reconstruct dynamic scenes from monocular depth estimation and optical flow. Robust-CVD proposes to first estimate the depth of each frame and then optimize camera poses and depth jointly [17]. CasualSAM leverages optical flow to train a depth and movement prediction network and optimize camera parameters [84]. MegaSaM proposes an optimization framework for dynamic scene reconstruction based on DROID-SLAM [21, 52]. Recently, DUSt3R [58] provides a powerful foundation model for static scene reconstruction. There are several recent extensions to DUSt3R. MonST3R fine-tunes DUSt3R on dynamic scene datasets to recover dynamic 3D scenes [82]. CUT3R introduces a continuous 3D perception model for online dense 3D reconstruction [57]. Easi3R identifies moving objects in the video from DUSt3R’s attention maps [3]. Aether proposes a unified model for reconstruction, planning, and prediction in dynamic scenes [50]. Our coarse prediction pipeline directly benefits from MonST3R [82], and our refinement framework is inspired by the optimization framework in CasualSAM [84] and Mega-SaM [21].
A key sub-task in dynamic scene understanding is segmenting the ‘moving map’ (mask image indicating which pixels are moving in the scene at each frame in the video). Recent approaches are mainly based on optical flow or point tracking [76, 32, 31, 68, 13, 8, 10]. However, we find that point tracking methods struggle with textureless objects and are sensitive to point locations (see supplement). Therefore, we do not use these approaches.
@@ 3.1 Problem Formulation
We define the interaction video as , where denotes the video length and each frame is an RGBD image with height and width . An object surface point cloud is assumed as input from a surface reconstruction stage prior to the start of interaction, as there are numerous methods for static surface reconstruction. The state of the object when reconstructing the surface is the initial state when recording the interaction video, and in practice can be captured in the same video right before the interaction begins.
We assume the video is casually captured with hand-held devices such as smartphones. Therefore, we cannot assume that the camera remains stationary during recording, nor do we assume that the camera trajectory can be easily obtained. We only assume the first video frame is aligned to the object surface coordinate system. In practice, this alignment can be achieved via point cloud registration or structure from motion. Normally, a human operator manipulates one joint at a time. Thus, we also assume only one part of the object is moving in the video. For multi-joint articulated objects, we can build each joint and part sequentially with consecutive interaction videos.
The output includes two key components: 1) joint parameters , including a binary value to indicate the type of joint (e.g., revolute or prismatic), as well as two triplets to indicate the position and orientation of the joint axis; and 2) movable parts of the object. Since it is challenging to segment movable parts on the 3D objects directly, we instead estimate a moving map for each video frame , which indicates the moving parts of the object in the video. Then, we unproject the moving map to 3D space with depth values to identify movable parts of the object. In addition, we also need to estimate both camera pose and the joint state at each frame.
@@ 3.2 Coarse Prediction
We first design a coarse prediction pipeline as shown in Fig.˜2. The idea is to decouple the problem by firstly finding the moving part of each frame of the video . Once we identify the dynamic region and static region in the video frames, we can use point motion in the static region to estimate the camera pose at each frame. We align all the observations to the same coordinate, and compute rigid transformations of movable parts of the object using points motion in the dynamic region to estimate joint parameters and joint state .
In practice, we use MonST3R [82] to provide a rough moving map of each video frame. Then we unproject pixels in to 3D point cloud with camera intrinsics, and use LoFTr [46] to compute pixel-level feature matching between two video frames. We use feature matching in static regions identified by to compute relative camera poses. We align all observations to the same camera coordinate system with the relative camera poses of each frame. Then, we use feature matching in the dynamic regions to compute a rigid transformation of the movable parts of the object. We can further estimate joint parameters from this rigid transformation. We estimate joint parameters from multiple video frame pairs and average the results as the coarse prediction of the joint parameters and states.
Through the coarse prediction stage, the proposed pipeline obtains rough results for moving map segmentation and joint parameters. However, this method has two disadvantages: 1) Since the moving map is inaccurate, camera poses and joint parameters estimation are inaccurate. 2) Feature matching provides a reliable prediction only when two video frames are visually similar. In other words, the joint states do not change much, and the point motion in dynamic regions will be minimal. In this situation, it’s hard to tell which joint types lead to this kind of point motion. Joint type prediction will be inaccurate as well.
@@ 3.3 Refinement
We refine the initial estimation results through gradient-based optimization. This phase is based on a simple idea: if the estimates of the joint parameters and camera poses are accurate, the object states observed in the video should transform back to the initial state. In other words, the transformed point cloud should align with the given object surface point cloud at the initial state. An overview of our pipeline is shown in Fig.˜3.
@@
Concretely, we divide the optimization target into static part and moving part. For the static part, we transform the point cloud to the same coordinate with camera poses . We compute a single-directional Chamfer distance to the given object surface point cloud in Eq.˜1.
@@
To compute the Chamfer distance of the moving part, we first compute the transformation matrix for the moving part with and . Then, we transform with camera pose and . Then, we compute the single-directional Chamfer distance to in Eq.˜2.
@@
Combining these two equations produces the final optimization objective . Note that this computation graph is differentiable. Thus, we can use gradient-based optimization methods to optimize joint parameters, camera poses, and the moving map.
@@
However, this naïve implementation has two problems. Firstly, it does not deal with newly observed parts of the object in the video. A newly observed part is a region that is initially hidden but becomes visible later (e.g., interior of a microwave). This region does not correspond with the object surface point cloud. Consequently, the Chamfer Distance is unreliable for point cloud alignment. Secondly, this implementation is quite inefficient. This is due to the need to optimize the probability of each pixel belonging to the moving part of each frame, which contains parameters in total. Considering each pixel individually is not efficient as there are millions of pixels in each frame.
To tackle the second problem our automatic part segmentation module aggregates pixels into object parts. This is more efficient and effective as the number of parts is much smaller than the number of pixels, reducing optimization complexity. Similar to Segment-Anything [16], automatic part segmentation segments the first video frame into several part regions, and then tracks these regions throughout the video using SAM2 [39]. Subsequently, instead of optimizing on each pixel, we optimize the probability of being a moving part for each tracked segment. We define a vector to represent the probability of each segment being a moving part, where is the number of segments. To construct for a frame, we multiply each probability value by the corresponding mask of the part. Finally, we sum the results across different parts to construct the moving map. We leverage an existing implementation of object part segmentation for videos.
The above automatic part segmentation module is also used to identify newly observed parts, thus helping to alleviate the first problem with the naïve implementation. If a part does not appear in the first frame, then it is classified as a newly observed part. By filtering out newly observed parts in the video, we reduce the spurious correspondence issue when computing the Chamfer Distance.
Since our pipeline only estimates the moving map of the input video, we need to find the movable part of the object on . We unproject the pixels in the moving region in the moving map of the first frame . Then, we select points on which are close to the unprotected moving pixels to be the movable part of the object. We use NKSR [9] to build meshes of both the static and the movable parts of the object.
| Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad or m) | Failure(%) |
| Revolute | Articulate Anything | 0.820.79 | 0.810.40 | 40.90 | N.A. | 38.63 |
| RSRD | 1.160.52 | 1.181.21 | 88.88 | 1.050.59 | 46.66 | |
| iTACO (Ours) | 0.320.56 | 0.130.25 | 15.90 | 0.250.46 | 6.81 | |
| iTACO w/o Refine | 1.030.67 | 0.240.25 | 68.18 | 0.590.47 | 4.54 | |
| iTACO w/o Coarse | 1.260.29 | 0.340.26 | 77.27 | 0.690.45 | 2.27 | |
| iTACO w/o Segment | 0.360.55 | 0.170.24 | 18.18 | 0.330.44 | 6.81 | |
| Prismatic | Articulate Anything | 0.920.78 | N.A. | 50.00 | N.A. | 46.66 |
| RSRD | 1.270.44 | N.A. | 40.00 | 0.630.41 | 36.66 | |
| iTACO | 0.240.33 | N.A. | 10.34 | 0.080.22 | 0 | |
| iTACO w/o Refine | 0.280.40 | N.A. | 6.89 | 0.160.16 | 0 | |
| iTACO w/o Coarse | 0.420.11 | N.A. | 6.89 | 0.080.09 | 0 | |
| iTACO w/o Segment | 0.300.41 | N.A. | 13.79 | 0.070.07 | 0 |